# & Télécom Paris, C2

### v5

@@ Abstract
Since its introduction in 1978, this is the first time an analysis111 we exclude results such as [28] that use an attack model for which direct countermeasures exist of the McEliece cryptosystem breaks the exponential barrier.
@@ 1 Introduction
Slightly more formally, a traditional security proof for the system relies on two assumptions:
@@
Although certain arguments, such as the fact that the class of Goppa codes is very large, make assumption 1 plausible, it remains quite ad hoc from a theoretical point of view. Its sole virtue is that it passed the test of time. There seems to be something special with Goppa codes happening there: indeed, variants of the McEliece system were proposed, with Goppa codes replaced with other types of codes allowing more manageable parameters; however, most of these propositions were eventually broken, as the hidden structure of the codes could be recovered. Also, although the McEliece cryptosystem marked the birth of code-based cryptography, the idea of having an object (such as a Goppa code) constructed from data defined over an extension field and then masked by considering it over a small subfield, was then found in other branches of cryptography: for instance it is at the basis of the HFE cryptosystem [24] in multivariate cryptography. Such systems can often be attacked by algebraic methods (including, but not limited to, the use of Gröbner basis algorithms). This suggests that if a weakness in assumption 1 were to be uncovered one day, then algebraic methods should be a tool of choice. However, up to now, the best distinguishers and key recovery algorithms such as those in [12][8][7][1] only apply to alternant or Goppa codes with very degraded parameters. Against McEliece with cryptographically relevant parameters, either they have exponential complexity with large constants, which makes them useless, or worse: they simply cease to work.
@@
Invalidating this belief, we present a new distinguisher for alternant and Goppa codes, i.e. a basic structural analysis of the McEliece cryptosystem, whose asymptotic complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover, for given finite parameters, it is more efficient than state-of-the-art distinguishers or key recovery methods, while not suffering from their strong regime limitations: in particular, it applies to the codes used in Classic McEliece.
@@
Likewise we claim that Theorem 2.8 of [10] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among $[7,4]$-codes,222erratum: among MDS $[7,4]$-codes (which then implies of full square) where we restrict to codes whose square is the whole space — otherwise we can use the square distinguisher. For such a code we always have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$. Let $Q_{1},Q_{2},Q_{3}$ be a basis of $I_{2}({\mathsf{C}})$. By definition, $Q_{1},Q_{2},Q_{3}$ do not satisfy linear relations with coefficients in $\mathbb{F}$, however they can satisfy relations whose coefficients are forms of degree $1$. Such relations are also called degree $3$ syzygies. They live in the kernel of the degree $3$ Macaulay matrix
@@
[…]
Thus we can distinguish by computing $\beta_{2,3}({\mathsf{C}})=\dim\ker(\mathbf{M}_{3})$, which will yield $2$ for $\operatorname{{\mathsf{GRS}}}$ and 00 for non-$\operatorname{{\mathsf{GRS}}}$ $[7,4]$-codes.
@@
Prior to that, section 2 explains how these invariants can be effectively computed. Last, section 5 combines everything and chooses parameters in order to optimize asymptotic complexity. Of special importance in this regard is Proposition 4, which shows that the property that the ideal of a code contains minors of a matrix of linear forms passes to its shortened subcodes.
@@
As already noted, our distinguisher can be seen as a generalization of the square distinguisher of [12]. Using an approach similar to ours, the work [7] also extends this square distinguisher by exploiting special properties of the space of quadratic relations, but in a different direction. Last, the key recovery attack in [1] combines shortening of the dual code with ideas from the square distinguisher, and a careful algebraic modeling in order to apply tools from Gröbner basis theory. All these results have limited range of applicability, but they introduced numerous techniques that influenced the present work.
@@
A $k$-dimensional subspace ${\mathsf{C}}\subseteq\mathbb{F}^{n}$ is called a $[n,k]$-code (and a $[n,k]_{q}$-code in case $\mathbb{F}=\mathbb{F}_{q}$). The orthogonal space ${\mathsf{C}}^{\perp}$ is called the dual code of ${\mathsf{C}}$.
@@ 2 Minimal resolutions and graded Betti numbers
Let $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the $k$-dimensional polynomial ring over $\mathbb{F}$, graded by total degree. If $M_{0}$ is a finitely generated graded $S$-module, and $F_{0}$ is the free module on a minimal system of homogeneous generators of $M_{0}$, then the (first) syzygy module of $M_{0}$ is $M_{1}=\ker(\;F_{0}\longrightarrow M_{0}\;)$. Concretely, if $g_{1},\dots,g_{N}$ are minimal generators of $M_{0}$, elements of $F_{0}$ can be seen as formal sums $\sum_{u}f_{u}[g_{u}]$ with $f_{u}\in S$, where the $[g_{u}]$ are just formal symbols; such a formal sum then lies in $M_{1}$ when the actual sum evaluated in $M_{0}$ satisfies $\sum_{u}f_{u}g_{u}=0$.
@@
where $S(-j)$ is the free rank $1$ module generated in degree $j$, so $S(-j)_{d}=S_{d-j}$.
@@
$\begin{array}[]{c|ccccc}&0&1&\dots&i&\dots\\ \hline\cr\vdots&\vdots&\vdots&&\vdots&\\ r&\beta_{0,r}&\beta_{1,r+1}&\dots&\beta_{i,r+i}&\dots\\ r+1&\beta_{0,r+1}&\beta_{1,r+2}&\dots&\beta_{i,r+i+1}&\dots\\ \vdots&\vdots&\vdots&&\vdots&\\ \end{array}$
with null entries marked as “--” for readability.
@@
From now on let ${\mathsf{C}}$ be a $[n,k]$-code, with generator matrix $\mathbf{G}$, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring. Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length $k-1$. So we have an exact sequence
@@
and the Betti diagram has $k$ columns, indexed from 00 to $k-1$.
As $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ and $F_{0}=S$, we have $\beta_{0,0}=1$ and $\beta_{0,j}=0$ for $j\neq 0$, i.e. the 00-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
@@
Hence the 00-th row of the Betti diagram is always $(1,-,-,\dots)$.
@@
$\begin{array}[]{c|cccc}&0&1&2&3\\ \hline\cr 0&1&-&-&-\\ 1&-&3&-&-\\ 2&-&1&6&3\\ \end{array}$
Figure 1: the $[7,4]_{2}$ Hamming code
$\begin{array}[]{c|cccccc}&0&1&2&3&4&5\\ \hline\cr 0&1&-&-&-&-&-\\ 1&-&10&16&-&-&-\\ 2&-&1&5&26&20&5\\ \end{array}$
Figure 2: the $[11,6]_{3}$ Golay code
$\begin{array}[]{c|cccccccccccc}&0&1&2&3&4&5&6&7&8&9&10&11\\ \hline\cr 0&1&-&-&-&-&-&-&-&-&-&-&-\\ 1&-&55&320&891&1408&1210&320&55&-&-&-&-\\ 2&-&1&11&55&220&650&1672&1870&1221&485&110&11\\ \end{array}$
Figure 3: the $[23,12]_{2}$ Golay code
@@
Effective computation
@@
[…]
@@
with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$ and columns indexed by $\mathcal{M}_{3}$: the row corresponding to $(X_{a},\mathbf{q})\in\mathcal{M}_{1}\times\mathcal{B}_{2}$, where $\mathbf{q}=(q_{M})_{M\in\mathcal{M}_{2}}$, has entry $q_{M}$ at position corresponding to $X_{a}M\in\mathcal{M}_{3}$, and 00 elsewhere.
@@
Now let $r\geq 4$, and assume inductively for all $13\leq i\leq r-$ 1 we have constructed a basis $\mathcal{B}_{i}$ of $M_{i-1,i}$, consisting of $\beta_{i-1,i}$ vectors, each of which has its entries indexed by $\mathcal{M}_{1}\times\mathcal{B}_{i-1}$.
@@
with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{r-1}$ and columns indexed by $\mathcal{M}_{2}\times\mathcal{B}_{r-2}$: the row corresponding to $(X_{a},\mathbf{s})\in\mathcal{M}_{1}\times\mathcal{B}_{r-1}$, where $\mathbf{s}=(s_{X_{b},\mathbf{t}})_{X_{b}\in\mathcal{M}_{1},\mathbf{t}\in\mathcal{B}_{r-2}}$, has entry $s_{X_{b},\mathbf{t}}$ at position corresponding to $(X_{a}X_{b},\mathbf{t})\in\mathcal{M}_{2}\times\mathcal{B}_{r-2}$, and 00 elsewhere.
@@
Observe that our computation only relies on mere linear algebra, and does not make use of any Gröbner basis theory. However, the two topics are clearly related. In particular, the many linear algebra optimizations used in Gröbner basis algorithms could certainly apply in our context.
@@
As the minimal resolution of a $[n,k]$-code ${\mathsf{C}}$ has length $k-1$, we always have $r_{\max}({\mathsf{C}})\leq k$.
@@ 3 Lower bounds from the Eagon-Northcott complex
Under mild hypotheses, this complex is exact [10, Th. A.2.60 & Th. 6.4], so it defines a resolution of the quotient $R/\operatorname{im}(\bigwedge^{g}\mathbf{\Phi})$ defined by the ideal generated by the maximal minors of $\mathbf{\Phi}$.
In this work we will only need the case $g=2$. In this case the complex has length $f-1$, and for $2\leq r\leq f$ its $(r-1)$-th module is free of rank
@@
$R[Z^{(j)}_{r;i_{1},\dots,i_{r}}]$ (52)
where, for each $2\leq r\leq f$, indices range over the $(r-1)\binom{f}{r}$ values $11\leq j\leq r-$ 1 and $1\leq i_{1}<\dots<i_{r}\leq f$, construct polynomials $s^{(j)}_{r;i_{1},\dots,i_{r}}$ as follows. First for $r=2$ we define the constant polynomials
$s^{(1)}_{2;i_{1},i_{2}}=x_{i_{1}}x^{\prime}_{i_{2}}-x^{\prime}_{i_{1}}x_{i_{2}}$ (53)
@@
$s^{(j)}_{r;i_{1},\dots,i_{r}}=\sum_{u=1}^{r}(-1)^{u-1}(x_{i_{u}}Z^{(j)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}}+x^{\prime}_{i_{u}}Z^{(j-1)}_{r-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}})$ (54)
where notation $\widehat{i_{u}}$ means $i_{u}$ is omitted, and we replace $Z^{(j)}_{r;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
@@
Either express (46) in the standard bases of $F=R^{f}$ and $G=R^{2}$. Or more directly, consider the $\mathbf{Z}_{r}$ and $\mathbf{s}_{r}$ as $(r-1)$-tuples of (exterior) $r$-vectors, observe that (53) means $\mathbf{s}^{(1)}_{2}=\mathbf{x}\wedge\mathbf{x}^{\prime}$, that (54) means $\mathbf{s}^{(j)}_{r}=\mathbf{x}\wedge\mathbf{Z}^{(j)}_{r-1}+\mathbf{x}^{\prime}\wedge\mathbf{Z}^{(j-1)}_{r-1}$, so the alternating property gives $\mathbf{s}^{(1)}_{3}(\mathbf{s}_{2})=\mathbf{x}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, $\;\mathbf{s}^{(2)}_{3}(\mathbf{s}_{2})=\mathbf{x}^{\prime}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, and for $r\geq 4$:
@@
So let ${\mathsf{C}}$ be a $[n,k]$-code over $\mathbb{F}$, and ${\mathsf{C}}_{\mathcal{H}}$ a codimension $1$ subcode. We will assume ${\mathsf{C}}$ projective, but then ${\mathsf{C}}_{\mathcal{H}}$ need not be so. Choose a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k-1}$ of ${\mathsf{C}}_{\mathcal{H}}$, and complete it to a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ of ${\mathsf{C}}$. Let $\mathbf{G}_{\mathcal{H}}$ and $\mathbf{G}$ be the corresponding generating matrices of ${\mathsf{C}}_{\mathcal{H}}$ and ${\mathsf{C}}$. Let also $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the polynomial ring in $k$ indeterminates, and $S_{\mathcal{H}}=\mathbb{F}[X_{1},\dots,X_{k-1}]$ its subring in the first $k-1$ indeterminates. We have a commutative diagram
@@
Let $f\geq g\geq d\geq 2$ be integers, and let $\mathbf{\Phi}\in S_{1}^{g\times f}$ be a matrix whose entries are homogeneous linear forms, i.e. elements of $S_{1}$. Recall from [10, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero. Let $V_{\mathbf{\Phi}}\subseteq S_{1}^{g}$ be the column span of $\mathbf{\Phi}$, so $\dim(V_{\mathbf{\Phi}})\leq f$, and set
@@
Application to alternant and Goppa codes
Consider a code ${\mathsf{C}}\in\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$, assumed to be proper. By (21), after extension of scalars, ${\mathsf{C}}_{\mathbb{F}_{q^{m}}}$ admits as a basis the $k=mt$ vectors $(\mathbf{y}\mathbf{x}^{a})^{q^{u}}$ for $10\leq a\leq t-$ 1 and $10\leq u\leq m-$ 1. Rename the $k=mt$ variables in our polynomial ring $S$ accordingly, so that our evaluation map now is
@@
The $2\times 2$ minor defined by the columns $\left[\begin{array}[]{c}X^{(e-u)}_{a}\\ X^{(e-u)}_{q^{u}+a}\end{array}\right]$ of $\mathbf{B}^{(e-u)}_{u}$ and $\left[\begin{array}[]{c}X^{(e-v)}_{b}\\ X^{(e-v)}_{q^{v}+b}\end{array}\right]$ of $\mathbf{B}^{(e-v)}_{v}$ is
$X^{(e-u)}_{a}X^{(e-v)}_{q^{v}+b}-X^{(e-u)}_{q^{u}+a}X^{(e-v)}_{b}$ (66)

### v6

@@ Abstract
Since its introduction in 1978, this is the first time an analysis (in the CPA model) of the McEliece cryptosystem breaks the exponential barrier.
@@ 1 Introduction
At the time of publication of [29] it was not customary to fully formalize the expected security properties of the scheme. In modern language we would say that it aims at one-wayness against chosen plaintext attacks (OW-CPA). There are now variants of the McEliece scheme that claim stronger security properties (e.g. IND-CCA2), but for this they rely on the hypothesis that the original McEliece scheme is OW-CPA. This makes it even more relevant today to scrutinize whether this hypothesis can be reduced to more fundamental theoretical assumptions.
And indeed, as the informal description of the system naturally suggests, a traditional security proof reduces this hypothesis to the following two assumptions:
@@
Assumption 1 was first introduced formally in [9]. Although certain arguments, such as the fact that the class of Goppa codes is very large, make this assumption plausible, it remains quite ad hoc from a theoretical point of view. Its sole virtue is that it passed the test of time. There seems to be something special with Goppa codes happening there: indeed, variants of the McEliece system were proposed, with Goppa codes replaced with other types of codes allowing more manageable parameters; however, most of these propositions were eventually broken, as the hidden structure of the codes could be recovered. Also, although the McEliece cryptosystem marked the birth of code-based cryptography, the idea of having an object (such as a Goppa code) constructed from data defined over an extension field and then masked by considering it over a small subfield, was then found in other branches of cryptography: for instance it is at the basis of the HFE cryptosystem [32] in multivariate cryptography. Such systems can often be attacked by algebraic methods (including, but not limited to, the use of Gröbner basis algorithms). This suggests that if a weakness in assumption 1 were to be uncovered one day, then algebraic methods should be a tool of choice. However, up to now, the best distinguishers and key recovery algorithms such as those in [15][11][10][2][24][23] only apply to alternant or Goppa codes with very degraded parameters. Against McEliece with cryptographically relevant parameters, either they have exponential complexity with large constants, which makes them useless, or worse: they simply cease to work.
@@
Preventing the spread of this belief, we present a new distinguisher for alternant and Goppa codes, i.e. a basic structural analysis of the McEliece cryptosystem, whose asymptotic complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover, for given finite parameters, it is more efficient than state-of-the-art distinguishers or key recovery methods, while not suffering from their strong regime limitations: in particular, it applies to the codes used in Classic McEliece.
@@
Likewise we claim that Theorem 2.8 of [13] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among MDS $[7,4]$-codes. Indeed, for such codes, the square always fills the whole space — so the square distinguisher does not apply — and we have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$. Let $Q_{1},Q_{2},Q_{3}$ be a basis of $I_{2}({\mathsf{C}})$. By definition, $Q_{1},Q_{2},Q_{3}$ do not satisfy linear relations with coefficients in $\mathbb{F}$, however they can satisfy relations whose coefficients are forms of degree $1$. Such relations are also called degree $3$ syzygies. They live in the kernel of the degree $3$ Macaulay matrix
@@
$X_{1}Q_{3}-X_{2}Q_{2}+X_{3}Q_{1}\;=\;X_{2}Q_{3}-X_{3}Q_{2}+X_{4}Q_{1}\;=\;0.$ (5)
Thus we can distinguish by computing $\beta_{2,3}({\mathsf{C}})=\dim\ker(\mathbf{M}_{3})$, which will yield $2$ for $\operatorname{{\mathsf{GRS}}}$ and 0 for non-$\operatorname{{\mathsf{GRS}}}$ MDS $[7,4]$-codes.
@@
Prior to that, section 2 explains how these invariants can be effectively computed. Last, section 5 combines everything and chooses parameters in order to optimize asymptotic complexity. Of special importance in this regard is Proposition 4 from section 3, which shows that the property that the ideal of a code contains minors of a matrix of linear forms passes to its shortened subcodes.
@@
As already noted, our distinguisher can be seen as a generalization of the square distinguisher of [15]. Using an approach similar to ours, the work [10] also extends this square distinguisher by exploiting special properties of the space of quadratic relations, but in a different direction: the authors detect the existence of short relations by computing the Hilbert series of a convenient Pfaffian modeling (and in some cases they even manage to recover the key). Last, the key recovery attack in [2] combines shortening of the dual code with ideas from the square distinguisher to reduce parameters, and then performs a careful algebraic modeling exploiting the symmetries of a system, that can finally be solved with Gröbner basis algorithms. All these results have limited range of applicability, but they introduced numerous techniques that influenced the present work.
After the first version of this text appeared, other works continued this series of results. A refined analysis of the Pfaffian distinguisher can now be found in [24]. Using purely geometric methods, [23] presents a new key recovery attack; but quite surprisingly, this attack works precisely in the regimes where those of [10] and [2] apply! Last, the recent work [41] proposes a distinguisher based on higher degree relations vanishing with higher multiplicities.
@@
Additional material and new results since the Eurocypt 2025 version
There are significative differences between this text and its prior version [37]. First, the supplementary material to [37] is now included in the main text, sometimes after some reworking. Moreover, several new results were introduced.
  1. (a)
    The discussion covering possible further improvements in the description of the minimal resolution of algebraic codes has been incorporated in section 3. After some reworking, it now includes Corollary 4 that provides an explicit family of linearly independent quadratic relations for dual alternant codes that matches the lower bound of [30, Th. 20] (thus giving an alternative proof for this result while sticking to the approach of [15]).
  2. (b)
    A tight lower bound on the number of columns of a matrix whose $2\times 2$ minors give quadratic relations (and determine the $r_{\max}$) of Goppa codes is now stated in Remark 1 in section 3 (although its proof is deferred to the forthcoming work [35]).
  3. (c)
    The computation of the Betti diagram of $[k,k+1]$ MDS codes and of [k,2k1][k,2k-1] $\operatorname{{\mathsf{GRS}}}$ codes has been incorporated in section 4.
  4. (d)
    The presentation of experimental data on defects of random codes also has been incorporated in section 4.
  5. (e)
    The new Proposition 12 in section 4 proves a (very small) part of Experimental fact 1, linking the $r_{\max}$ of a code to its minimum distance.
  6. (f)
    The two examples of application of the distinguisher to some Goppa codes with small parameters have been incorporated in section 5.
  7. (g)
    A new variant of the distinguisher, based on Koszul cohomology and the (block) Wiedemann algorithm, has been introduced also in section 5.
Some of these new results allow a noticeable improvement on the complexity of the distinguisher. Point (b) allows to predict the optimal number $s$ of coordinates to shorten. Point (g) allows to use linear algebra with exponent $2$ instead of $\omega$ (and also to relax the dependency of the complexity estimate from some of the heuristics). When applied e.g. to the Classic McEliece 348864 parameters, each allows to gain roughly 60 bits of complexity, and combined, they reduce the complexity estimate from $2^{528}$ to $2^{401}$.
@@
A $k$-dimensional subspace ${\mathsf{C}}\subseteq\mathbb{F}^{n}$ is called a $[n,k]$-code (and a $n,k]_{q}-$code in case $\mathbb{F}=\mathbb{F}_{q}$). The orthogonal space $\mathsf{C}$$^{\perp}$ is called the dual code of $\mathsf{C}$}.
@@ 2 Minimal resolutions and graded Betti numbers
Let $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the $k$-dimensional polynomial ring over $\mathbb{F}$, graded by total degree. If $M_{0}$ is a finitely generated graded $S$-module, and $F_{0}$ is the free module on a minimal system of homogeneous generators of $M_{0}$, then the (first) syzygy module of $M_{0}$ is $M_{1}=\ker(\;F_{0}\longrightarrow M_{0}\;)$. Concretely, if $g_{1},\dots,g_{N}$ are minimal generators of $M_{0}$, elements of $F_{0}$ can be seen as formal sums $\sum_{u}f_{u}[g_{u}]$ with $f_{u}\in S$, where the $g_{u}]$ are just formal symbols; such a formal sum then lies in $M_{1}$ when the actual sum evaluated in $M_{0}$ satisfies $\sum_{u}f_{u}g_{u}=0$.
@@
where S(j)S(-j) is the free rank $1$ module generated in degree $j$, so $S(-j)_{d}=S_{d-j}$.
@@
[…]
with null entries marked as “-” for readability.
@@
From now on let $\mathsf{C}$} be a $[n,k]$-code, with generator matrix $\mathbf{G}$, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring. Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length k1k-1. So we have an exact sequence
@@
and the Betti diagram has $k$ columns, indexed from 0 to k1k-1.
As $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ and $F_{0}=S$, we have $\beta_{0,0}=1$ and $\beta_{0,j}=0$ for $j\neq 0$, i.e. the 0-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
@@
Hence the 0-th row of the Betti diagram is always $(1,-,-,\dots)$.
@@
[…]
Figure 1: the $7,4]_{2}$ Hamming code
[…]
Figure 2: the $11,6]_{3}$ Golay code
[…]
Figure 3: the $23,12]_{2}$ Golay code
@@
Effective computation of linear syzygies
@@
$M_{r-2,r}$$M_{r-2,r-1}\otimes S_{1}$$F_{r-3,r}$$M_{r-3,r-2}\otimes S_{2}=F_{r-3,r-2}\otimes S_{2}$$\scriptstyle{\varphi_{r}}$$\scriptstyle{\psi_{r}}$ (32)
@@
with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{2}$ and columns indexed by $\mathcal{M}_{3}$: the row corresponding to $(X_{a},\mathbf{q})\in\mathcal{M}_{1}\times\mathcal{B}_{2}$, where $\mathbf{q}=(q_{M})_{M\in\mathcal{M}_{2}}$, has entry $q_{M}$ at position corresponding to $X_{a}M\in\mathcal{M}_{3}$, and 0 elsewhere.
@@
Now let $r\geq 4$, and assume inductively for all $3\leq i\leq r-1$ we have constructed a basis $\mathcal{B}_{i}$ of $M_{i-1,i}$, consisting of $\beta_{i-1,i}$ vectors, each of which has its entries indexed by $\mathcal{M}_{1}\times\mathcal{B}_{i-1}$.
@@
with rows indexed by $\mathcal{M}_{1}\times\mathcal{B}_{r-1}$ and columns indexed by $\mathcal{M}_{2}\times\mathcal{B}_{r-2}$: the row corresponding to $(X_{a},\mathbf{s})\in\mathcal{M}_{1}\times\mathcal{B}_{r-1}$, where $\mathbf{s}=(s_{X_{b},\mathbf{t}})_{X_{b}\in\mathcal{M}_{1},\mathbf{t}\in\mathcal{B}_{r-2}}$, has entry $s_{X_{b},\mathbf{t}}$ at position corresponding to $(X_{a}X_{b},\mathbf{t})\in\mathcal{M}_{2}\times\mathcal{B}_{r-2}$, and 0 elsewhere.
@@
An implementation can be found in [34], which was used to compute all the examples in this text.
Observe that this algorithm only relies on mere linear algebra, and does not make use of any Gröbner basis theory. However, the two topics are clearly related. In particular, the many linear algebra optimizations used in Gröbner basis algorithms certainly apply here also. This will be discussed later in the text, when we’ll deal with complexity estimates.
@@
As the minimal resolution of a $[n,k]$-code $\mathsf{C}$} has length k1k-1, we always have $r_{\max}({\mathsf{C}})\leq k$.
@@ 3 Lower bounds from the Eagon-Northcott complex
Under mild hypotheses, this complex is exact [13, Th. A.2.60 & Th. 6.4], so it defines a resolution of the quotient $R/\operatorname{im}(\bigwedge^{g}\mathbf{\Phi})$ defined by the ideal generated by the maximal minors of $\mathbf{\Phi}$.
In this work we will only need the case $g=2$. In this case the complex has length f1f-1, and for $2\leq r\leq f$ its (r1)(r-1)-th module is free of rank
@@
$R[Z_{r,j;i_{1},\dots,i_{r}}]$ (52)
where, for each $2\leq r\leq f$, indices range over the $(r-1)\binom{f}{r}$ values $1\leq j\leq r-1$ and $1\leq i_{1}<\dots<i_{r}\leq f$, construct polynomials $s_{r,j;i_{1},\dots,i_{r}}$ as follows. First for $r=2$ we define the constant polynomials
$s_{2,1;i_{1},i_{2}}=x_{i_{1}}x^{\prime}_{i_{2}}-x^{\prime}_{i_{1}}x_{i_{2}}$ (53)
@@
$s_{r,j;i_{1},\dots,i_{r}}=\sum_{u=1}^{r}(-1)^{u-1}(x_{i_{u}}Z_{r-1,j;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}}+x^{\prime}_{i_{u}}Z_{r-1,j-1;i_{1},\dots,\widehat{i_{u}},\dots,i_{r}})$ (54)
where notation $\widehat{i_{u}}$ means $i_{u}$ is omitted, and we replace $Z_{r,j;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
@@
Either express (46) in the standard bases of $F=R^{f}$ and $G=R^{2}$. Or more directly, consider the $\mathbf{Z}_{r}$ and $\mathbf{s}_{r}$ as (r1)(r-1)-tuples of (exterior) $r$-vectors, observe that (53) means $\mathbf{s}_{2,1}=\mathbf{x}\wedge\mathbf{x}^{\prime}$, that (54) means $\mathbf{s}_{r,j}=\mathbf{x}\wedge\mathbf{Z}_{r-1,j}+\mathbf{x}^{\prime}\wedge\mathbf{Z}_{r-1,j-1}$, so the alternating property gives $\mathbf{s}_{3,1}(\mathbf{s}_{2})=\mathbf{x}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, $\mathbf{s}_{3,2}(\mathbf{s}_{2})=\mathbf{x}^{\prime}\wedge\mathbf{x}\wedge\mathbf{x}^{\prime}=\mathbf{0}$, and for $r\geq 4$:
@@
So let $\mathsf{C}$} be a $[n,k]$-code over $\mathbb{F}$, and $\mathsf{C}$$_{\mathcal{H}}$ a codimension $1$ subcode. We will assume $\mathsf{C}$} projective, but then $\mathsf{C}$$_{\mathcal{H}}$ need not be so. Choose a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k-1}$ of $\mathsf{C}$$_{\mathcal{H}}$, and complete it to a basis $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$ of $\mathsf{C}$}. Let $\mathbf{G}_{\mathcal{H}}$ and $\mathbf{G}$ be the corresponding generating matrices of $\mathsf{C}$$_{\mathcal{H}}$ and $\mathsf{C}$}. Let also $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the polynomial ring in $k$ indeterminates, and $S_{\mathcal{H}}=\mathbb{F}[X_{1},\dots,X_{k-1}]$ its subring in the first k1k-1 indeterminates. We have a commutative diagram
@@
Let $f\geq g\geq d\geq 2$ be integers, and let $\mathbf{\Phi}\in S_{1}^{g\times f}$ be a matrix whose entries are homogeneous linear forms, i.e. elements of $S_{1}$. Recall from [13, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero. Let $V_{\mathbf{\Phi}}\subseteq S_{1}^{g}$ be the column span of $\mathbf{\Phi}$, so $\dim(V_{\mathbf{\Phi}})\leq f$, and set
@@
Application to alternant codes
Consider a code ${\mathsf{C}}\in\operatorname{{\mathsf{Alt}}}^{\perp}_{q,m,n,t}$, assumed to be proper. By (21), after extension of scalars, ${\mathsf{C}}_{\mathbb{F}_{q^{m}}}$ admits as a basis the $k=mt$ vectors $(\mathbf{y}\mathbf{x}^{a})^{q^{u}}$ for $0\leq a\leq t-1$ and $0\leq u\leq m-1$. Rename the $k=mt$ variables in our polynomial ring $S$ accordingly, so that our evaluation map now is
@@
The $2\times 2$ minor defined by the columns $\left[\begin{array}[]{c}X^{(e-u)}_{a}\\ X^{(e-u)}_{a+q^{u}}\end{array}\right$] of $\mathbf{B}^{(e-u)}_{u}$ and $\left[\begin{array}[]{c}X^{(e-v)}_{b}\\ X^{(e-v)}_{b+q^{v}}\end{array}\right$] of $\mathbf{B}^{(e-v)}_{v}$ is
$X^{(e-u)}_{a}X^{(e-v)}_{b+q^{v}}-X^{(e-v)}_{b}X^{(e-u)}_{a+q^{u}}$ (66)